OpenStudy (vishweshshrimali5):

\[\text {Let,}~~ \text{x} = x - [x]~.~ \text{If }~{\{a}^{-1}\} = {\{a^2}\}~ \text {and}~ 2 < a^2 < 3,~ \text{then} a^{12} - 144*a^{-1} = ?\]

OpenStudy (vishweshshrimali5):

I know that, \([a^2] = 2\)

OpenStudy (vishweshshrimali5):

Sorry couldn't get the correct latex.

mathslover (mathslover):

\[\text {Let,}~~ \text{x} = x - [x]~.~ \text{If }~{{a}^{-1}} = {{a^2}}~ \text {and}~ 2 < a^2 < 3,~ \text{then} a^{12} - 144*a^{-1} = ?\]

OpenStudy (vishweshshrimali5):

{a^{-1} = a^{2}} THis should be present

OpenStudy (vishweshshrimali5):

\(\text {Let,}~~ \text{x} = x - [x]~.~ \text{If }~{\{a}^{-1}\} = {\{a^2}\}~ \text {and}~ 2 < a^2 < 3,~ \text{then} a^{12} - 144*a^{-1} = ?\)

OpenStudy (vishweshshrimali5):

This is the correct question.

OpenStudy (rational):

\(\text {Let,}~~ \text{{x}} = x - [x]~.~ \\ \text{If }~{\{a}^{-1}\} = {\{a^2}\}~ \text {and}~ 2 < a^2 < 3,~ \text{then } a^{12} - 144*a^{-1} = ?\)

OpenStudy (rational):

I think you mean that ?

OpenStudy (vishweshshrimali5):

yes

OpenStudy (isaiah.feynman):

I got two "solutions" to that...

OpenStudy (rational):

[a] = 1 [1/a] = 0 {a^{-1}} = 1/a {a^2} = a^2-1

OpenStudy (rational):

1/a = a^2 - 1

OpenStudy (isaiah.feynman):

Is this the golden number???

OpenStudy (rational):

its a cubic^

OpenStudy (isaiah.feynman):



OpenStudy (isaiah.feynman):